# Diffusion differentiable resampling

### v1

@@ Abstract
This paper is concerned with differentiable resampling in the context of sequential Monte Carlo (e.g., particle filtering). We propose a new informative resampling method that is instantly pathwise differentiable, based on an ensemble score diffusion model. We prove that our diffusion resampling method provides a consistent estimate to the resampling distribution, and we show by experiments that it outperforms the state-of-the-art differentiable resampling methods when used for stochastic filtering and parameter estimation.
@@ 1 Introduction
Consider a distribution $\pi$ and a population of weighted samples $\{(w_{i},X_{i})\}_{i=1}^{N}\sim\pi$, where $\{X_{i}\}_{i=1}^{N}$ are often identically and independently drawn from another proposal distribution. The goal of resampling is to transform these weighted samples into an un-weighted set while preserving the original distribution $\pi$. In the weak sense, unbiased resampling is defined as a mapping
@@
for any bounded and continuous test function $\psi$, where $\{X_{i}^{*}\}_{i=1}^{N}$ stands for the re-samples.
@@
Resampling is a key component in sequential Monte Carlo (SMC) samplers for state-space models (SSMs, ChopinBook2020). It not only mitigates particle degeneracy in practice, but has also been shown to represent a jump Markov process in a continuous-time limit (Chopin2022PDMP). However, resampling typically hinders (gradient-based) parameter estimation of SSMs due to discrete randomness. For instance, with the commonly used multinomial resampling, one draws indices $I_{i}\sim\mathrm{Categorical}(w_{1},w_{2},\ldots,w_{N})$ independently for $i=1,2,\ldots,N$, and then define the resampling mapping by indexing $X^{*}_{i}\coloneqq X_{I_{i}}$. If the sample $X^{\theta}_{i}$ (and weight) depend on some unknown parameters $\theta$, the pathwise derivative of the re-sample, $\partial X_{i}^{\theta,*}/\,\partial\theta$, is not defined. Moreover, most automatic differentiation libraries (e.g., JAX) will typically drop the undefined derivatives, resulting in erroneous gradient estimates (see, e.g., Naesseth2018).
@@
Various methods have thus been proposed to make the resampling step differentiable in the context of SMC and SSMs (see recent surveys in Xiongjie2025; Brady2025pydpf). One line of work focuses on the expectation derivative $\partial\operatorname{\mathbb{E}}\bigl[X_{i}^{\theta,*}\bigr]/\,\partial\theta$, which is often well defined, rather than the pathwise derivative. This can be achieved by combing Fisher’s score (Poyiadjis2011) and stochastic derivatives (Arya2022), resulting in, for example, a stop-gradient based method by Scibior2021differentiable. However, these methods often suffer from high variance and may consequently require a large sample size $N$.
@@
Another line of work focuses on developing new resampling methods that naturally comes with well-defined pathwise derivatives, in contrast to multinomial and its variants (e.g., residual resampling). Notable examples include soft (Karkus2018) and Gumbel-Softmax resampling (Jang2017categorical). Both methods essentially form an interpolation between multinomial resampling (which has no gradient) and uninformative resampling (which has a gradient) via a calibration parameter. Although they have been empirically shown to work well for certain models, they are fundamentally biased. Crucially, one has to decide on a trade-off between the gradient bias and the statistical performance of the resampling mapping. Similarly, Zhu2020towards propose to parametrise the resampling with a neural network, which adds both training complexity and additional sources of bias.
@@
The perhaps first fully-differentiable-by-construction resampling method with consistency guarantees is due to Malik2011, where they make a smooth approximation to the empirical cumulative distribution function (CDF) of the weighted samples. However, their method is applicable to univariate $X$ only, and this was later generalised by Li2024differentiable using kernel jittering to approximate the CDF gradient. In a similar vein, Corenflos2021 propose an optimal transport (OT) based resampling method. The idea is to learn a (linear) transportation map between the target distribution $\pi$ and the proposal, and approximate the resampling by an ensemble transformation (Reich2013)
@@
where $P^{\varepsilon}\in\mathbb{R}^{N\times N}$ denotes the $\varepsilon$-regularised entropic optimal coupling. The main problem of OT-based resampling lies in computation, that one has to compute for the transportation plan $P^{\varepsilon}$. The cost scales quadratically in the number of samples $N$ (when implemented Sinkhorn naïvely) which in turn depends exponentially on the entropy parameter $1\,/\,\varepsilon$ (Luo2023OT; Burns2025linear) to converge. In the context of SMC, the method may not work well when the proposal/reference does not well approximate the target. Li2024differentiable also show a (hypothetical) case when linear transformation of OT is insufficient for exploiting the distribution manifold. Nevertheless, this line of work has inspired us to develop a new transportation-based resampling method that avoids these issues.
Other than making the resampling differentiable, one can also modify the SMC algorithm itself to produce smooth estimates of the marginal log likelihood of SSMs. For instance, Klaas2005 introduce a mixture of SMC proposals to marginalise out the need for resampling, however, in practice, drawing samples from mixture distributions typically requires discrete sampling too. This was addressed by Lai2022 using implicit reparametrisation, but the approach remains restricted to structured proposals. Overall, this class of methods rely on customised SMC samplers, potentially limiting their applicability in general.
Therefore, our main motivation in this paper is to develop a differentiable resampling method that can be generically applied as is, without altering the SMC (or SSMs) construction, sacrificing consistency, or increasing computational cost. We take inspiration from the transportation-based approach (Corenflos2021), but our key departure here is the transport map construction: it needs not to be computed but rather specified, thereby mitigating the computation issue. Moreover, our construction allows for integrating additional information of the target into the map to make it statistically more efficient. Our contributions are as follows.
  • We introduce diffusion resampling, a new resampling paradigm that instantly allows for automatic differentiation $\partial X_{i}^{\theta,*}/\,\partial\theta$ and consequently also for expectation $\partial\operatorname{\mathbb{E}}\bigl[X_{i}^{\theta,*}\bigr]/\,\partial\theta$. We apply the method for filtering and gradient-based parameter estimation in state-space models with SMC samplers.
  • We prove that our diffusion resampling method is consistent in the number of samples. We shown an informative error bound in Wasserstein distance explicitly quantifying the error propagation of the resampling.
@@ 2 Diffusion differentiable resampling
In this section we show how we can make use of a diffusion model (without training, cf. Baker2025; Wan2025diffpf) to construct a differentiable resampling scheme and apply it for sequential Monte Carlo. The idea is akin to Equation (2), which computes a linear transportation map, but here we instead use a diffusion model to construct a non-linear map. We define the diffusion model via a (forward-time) Langevin stochastic differential equation (SDE)
@@
initialised at the target $\pi$, where $\pi_{\mathrm{ref}}$ is a user-chosen reference distribution from which we can easily sample (e.g., Gaussian), $W$ is a Brownian motion, and $b$ is a dispersion constant. Under mild conditions (Meyn2009), the marginal distribution $p_{t}$ of $X(t)$ converges to $\pi_{\mathrm{ref}}$ geometrically fast as $t\to\infty$. Importantly, Song2021scorebased; Anderson1982 show that we can leverage this construction to sample from $\pi$ if we can sample $p_{T}$ at some terminal time $T>0$ and simulate the reverse-time SDE [^1] To simplify later analysis we assume that the Brownian motions in Equations (3) and (4) are the same.
@@
At time $T$, the marginal distribution $q_{T}$ of $U(T)$ equals $\pi$ by construction, since X(Tt)X(T-t) and $U(t)$ solve the same Kolmogorov forward equation. Resampling can thus be seen as sampling from this reversal at $T$. The challenge is that the score function $\operatorname{\nabla\!}\log p_{t}$ is intractable, and in the context of generative diffusion models the score is usually learnt from samples of $\pi$, introducing demanding computations. The same issue also happens for other constructions of the forward-reverse pairs, such as dynamic Schrödinger bridges (DeBortoli2021diffusion) and stochastic interpolants (Albergo2023stochastic); see Zhao2024rsta for a review. However, given that we have access to $\{(w_{i},X_{i})\}_{i=1}^{N}\sim\pi$, we can approximate the score (BaoEnsScore2024) without training according to
@@
where $\alpha_{i}(x,t)\coloneqq w_{i}\,p_{t|0}(x{\;|\;}X_{i})\,/\sum_{j=1}^{N}w_{j}\,p_{t|0}(x{\;|\;}X_{j})$ stands for the normalised ensemble weight. This approximation exactly functions as an importance sampling, where $\pi(x_{0})$ and $p_{t|0}(\cdot{\;|\;}x_{0})$ stand for the prior/proposal and likelihood, respectively. As such, the established $N\to\infty$ consistency properties of (auto-normalised) importance sampling apply (ChopinBook2020) at least pointwise for $(x,t)\mapsto s_{N}(x,t)$. Evaluation of the function $s$ has an $O(N)$ computational cost if implemented naïvely, and an logarithmic cost if implemented in parallel (Lee01012010).
Therefore, the resampling can be approximately achieved by simulating the reverse SDE (4) using the ensemble score $s_{N}$ in Equation (6) until a terminal time $T$. Similar to Corenflos2021, this diffusion process too defines an optimal transportation but in the sense of Jordan–Kinderlehrer–Otto scheme (Jordan1998). The key distinction is that this map is given by construction, and does not need to be computed like in OT with Sinkhorn. Although the diffusion also assumes $T\to\infty$, we show in Section 3 that $T$ scales better than the entropy parameter $1\,/\,\varepsilon$ as a function of $N$.
@@ 2.1 Differentiable sequential Monte Carlo
The diffusion resampling in Algorithm 1 is particularly useful for SMC and SSMs for two reasons. First, the resampling method is pathwise-differentiable by construction. Moreover, recent advances provide methods for propagating gradients through SDE solvers (see, e.g., Bartosh2025sde; LiXueChen2020; Kidger2021gradient) and these methods have been well implemented in common automatic differentiation libraries (Kidger2021on). Secondly, we can fully leverage the sequential structure of SMC to adaptively choose for a suitable reference distribution $\pi_{\mathrm{ref}}$ based on any previous SMC marginal distribution. To see these aspects, let us begin by considering a parametrised Feynman–Kac model
$\begin{split}Q^{\theta}_{0:J}(z_{0:J})&=\frac{1}{L(\theta)}\,M_{0}^{\theta}(z_{0})\,G_{0}^{\theta}(z_{0})\\ &\quad\times\prod_{j=1}^{J}M_{j}^{\theta}(z_{j}{\;|\;}z_{j-1})\,G_{j}^{\theta}(z_{j},z_{j-1}),\end{split}$ (8)
where $M^{\theta}_{j}$ and $G_{j}^{\theta}$ are Markov transition and potential functions, respectively, and $L(\theta)$ is the marginal likelihood that we often aim to maximise. Take an SSM with state transition $p_{\theta}(z_{j}{\;|\;}z_{j-1})$ and measurement $p_{\theta}(y_{j}{\;|\;}z_{j})$ for example. In this case a bootstrap construction of the corresponding Feynman–Kac model is simply $M_{j}^{\theta}(z_{j}{\;|\;}z_{j-1})=p_{\theta}(z_{j}{\;|\;}z_{j-1})$ and $G_{j}^{\theta}(z_{j},\cdot)=p_{\theta}(y_{j}{\;|\;}z_{j})$. One can draw samples of a Feynman–Kac model with an SMC sampler as in Algorithm 2. For detailed exposition of Feynman–Kac models and SMC samplers, we refer the readers to ChopinBook2020; DelMoral2004.
@@ 16   Weight $w_{j,i}\leftarrow\overline{w}_{j,i}\,/\,L_{j}(\theta)$
This choice of the reference $\pi_{\mathrm{ref}}$ is more flexible compared to that of Corenflos2021 who explicitly used the predictive samples $\{(w_{j-1,i},Z_{j,i})\}_{i=1}^{N}$ at the $j$-th SMC step as the reference to resample $\{(w_{j,i},Z_{j,i})\}_{i=1}^{N}$. In contrast, diffusion resampling allows to use, e.g., the posterior samples $\{(w_{j,i},Z_{j,i})\}_{i=1}^{N}$ to establish the reference, which can be more informative than the predictive one.
@@ 2.2 Mean-reverting Gaussian reference
A remaining question is how to choose the reference distribution $\pi_{\mathrm{ref}}$. In generative sampling, the reference is usually a unit Normal. However, this becomes suboptimal for resampling when the target distribution $\pi$ is geometrically far away from $\mathrm{N}(0,I_{d})$, and as a consequence we would need large enough $T$ for convergence. A more informative choice of $\pi_{\mathrm{ref}}$ is a Gaussian approximation of $\pi$. Given that we have access to the target samples $\{(w_{i},X_{i})\}_{i=1}^{N}\sim\pi$, we can make use of moment matching, where $\mu_{N}\coloneqq\sum_{i=1}^{N}w_{i}\,X_{i}$ and $\Sigma_{N}\coloneqq\sum_{i=1}^{N}w_{i}\,(X_{i}-\mu_{N})\,(X_{i}-\mu_{N})^{\mkern-1.5mu\mathsf{T}}$ stand for the empirical mean and covariance, respectively (Yang2013; Kang2025). We can then choose the reference measure to be
@@
$\begin{split}p_{t|0}(x_{t}{\;|\;}x_{0})&=\mathrm{N}(x_{t};m_{t}(x_{0}),V_{t}),\\ m_{t}(x_{0})&\coloneqq x_{0}\,\mathrm{e}^{-b^{2}\,\Sigma_{N}^{-1}\,t}+\mu_{N}\,\bigl(1-\mathrm{e}^{-b^{2}\,\Sigma_{N}^{-1}\,t}\bigr),\\ V_{t}&\coloneqq\Sigma_{N}^{-1}\,\bigl(1-\mathrm{e}^{-2\,b^{2}\,\Sigma_{N}^{-1}\,t}\bigr).\end{split}$
To combine with the SMC in Algorithm 2, we compute $\mu_{N}$ and $\Sigma_{N}$ based on $\bigl\{(w_{j-1,i},Z_{j-1,i})\bigr\}_{i=1}^{N}$, see Appendix A for details. It is also possible to leverage any Gaussian filter which is commonly used for approximating the optimal proposal (Vandermerwe2000), to establish the reference. The central gist here is to exploit the sequential structure of SMC to adaptively and informatively choose the reference, instead of assuming a static and uninformative one, such as $\mathrm{N}(0,I_{d})$. Similar ideas of using mean-reverting SDEs for informative generative sampling has been used in domain-specific applications, such as image restoration (Luo2023IRSDE; Luo2024rsta).
In practice, to avoid solving the inversion $\Sigma_{N}^{-1}$, we can approximate it as a diagonal, or directly estimate an empirical precision matrix (Yuan2007; Fan2016). When the Gaussian construction is insufficient for multi-mode targets, one may use a Gaussian mixture, although the associated semigroup needs to be approximated. Another option is to transform with a diffeomorphism to obtain a flexible yet tractable reference process (Deng2020).

### v2

@@ Abstract
This paper is concerned with differentiable resampling in the context of sequential Monte Carlo (e.g., particle filtering). We propose a new informative resampling method that is instantly differentiable, based on an ensemble score diffusion model. We theoretically prove that our diffusion resampling method provides a consistent resampling distribution, and we show empirically that it outperforms the state-of-the-art differentiable resampling methods on multiple filtering and parameter estimation benchmarks. Finally, we show that it achieves competitive end-to-end performance when used in learning a complex dynamics-decoder model with high-dimensional image observations.
@@ 1 Introduction
Consider a distribution $\pi$ and a population of weighted samples $\{(w_{i},X_{i})\}_{i=1}^{N}\sim\pi$, where $\{X_{i}\}_{i=1}^{N}$ are often identically and independently drawn from another proposal distribution, calibrated by the weights. The goal of resampling is to transform these weighted samples into an un-weighted set while preserving the original distribution $\pi$. In the weak sense, unbiased resampling is defined as a mapping
@@
for any bounded and continuous test function $\psi$, where $\{X_{i}^{*}\}_{i=1}^{N}$ stands for the re-samples, and $X\sim\pi$.
@@
Resampling is a key component in sequential Monte Carlo (SMC) samplers for state-space models (SSMs, Chopin and Papaspiliopoulos, 2020). It not only mitigates particle degeneracy in practice, but has also been shown to represent a jump Markov process in a continuous-time limit (Chopin et al., 2022). However, resampling typically hinders (gradient-based) parameter estimation of SSMs due to discrete randomness. For instance, with the commonly used multinomial resampling, one draws indices $I_{i}\sim\mathrm{Categorical}(w_{1},w_{2},\ldots,w_{N})$ independently for $i=1,2,\ldots,N$, and then define the resampling mapping by indexing $X^{*}_{i}\coloneqq X_{I_{i}}$. If the sample $X^{\theta}_{i}$ (and weight) depend on some unknown parameters $\theta$, the pathwise derivative of the re-sample, $\partial X_{i}^{\theta,*}/\,\partial\theta$, is not defined. Moreover, most automatic differentiation libraries (e.g., JAX) will typically drop the undefined derivatives, resulting in erroneous gradient estimates (see, e.g., Naesseth et al., 2018).
@@
Various methods have thus been proposed to make the resampling step differentiable in the context of SMC and SSMs (see recent surveys in Chen and Li, 2025; Brady et al., 2025). One line of work focuses on the expectation derivative $\partial\operatorname{\mathbb{E}}\bigl[X_{i}^{\theta,*}\bigr]/\,\partial\theta$, which is often well defined, rather than the pathwise derivative. This can be achieved by combing Fisher’s score (Poyiadjis et al., 2011) and stochastic derivatives (Arya et al., 2022), resulting in, for example, a stop-gradient based method by Ścibior and Wood (2021). However, these methods often suffer from high variance and may consequently require a large sample size $N$.
@@
Another line of work focuses on developing new resampling methods that naturally come with well-defined pathwise derivatives, in contrast to multinomial and its variants (e.g., residual resampling). Notable examples include soft (Karkus et al., 2018) and Gumbel-Softmax resampling (Jang et al., 2017). Both methods essentially form an interpolation between multinomial resampling (which has no gradient) and uninformative resampling (which has a gradient) via a calibration parameter. Although they have been empirically shown to work well for certain models, they are fundamentally biased. Crucially, one has to decide on a trade-off between the gradient bias and the statistical performance of the resampling mapping. Similarly, Zhu et al. (2020) propose to parametrise the resampling with a neural network, which adds both training complexity and additional sources of bias.
@@
The perhaps first fully-differentiable-by-construction resampling method with consistency guarantees is due to Malik and Pitt (2011), who make a smooth approximation of the empirical cumulative distribution function (CDF) of the weighted samples. However, their method is applicable to univariate $X$ only. This was later generalised by Li et al. (2024) using kernel jittering to approximate the CDF gradient. In a similar vein, Corenflos et al. (2021) propose an optimal transport (OT) based resampling method. The idea is to learn a (linear) transportation map between the target distribution $\pi$ and the proposal, and approximate the resampling by an ensemble transformation (Reich, 2013)
@@
where $P^{\varepsilon}\in\mathbb{R}^{N\times N}$ denotes the $\varepsilon$-regularised entropic optimal coupling. The main problem of OT-based resampling lies in computation, that one has to compute for the transportation plan $P^{\varepsilon}$. The cost scales quadratically in the number of samples $N$ with aSinkhorn implementation, which in turn depends exponentially on the entropy parameter $1\,/\,\varepsilon$ (Luo et al., 2023a; Burns and Liang, 2025) to converge. In the context of SMC, the method may not work well when the proposal/reference does not well approximate the target. Li et al. (2024, Fig. 1) also show a (hypothetical) case when linear transformation of OT is insufficient for exploiting the distribution manifold. Nevertheless, this line of work has inspired us to develop a new transportation-based resampling method that avoids these issues.
Other than making the resampling differentiable, one can also modify the SMC algorithm itself to produce smooth estimates of the marginal log likelihood of SSMs. For instance, Klaas et al. (2005) introduce a mixture of SMC proposals to marginalise out the need for resampling, however, in practice, drawing samples from mixture distributions typically requires discrete sampling too. This was addressed by Lai et al. (2022) using implicit reparametrisation, but the approach remains restricted to structured proposals. Overall, this class of methods rely on customised SMC samplers, potentially limiting their applicability in general.
Therefore, our main motivation in this paper is to develop a differentiable resampling method that can be generically applied as is, without altering the SMC (or SSMs) construction, sacrificing consistency, or increasing computational cost. We take inspiration from the transportation-based approach (Corenflos et al., 2021), but our key departure here is the transport map construction: it needs not to be computed but rather specified, thereby mitigating the computation issue. Moreover, our construction allows for integrating additional information of the target into the map to make it statistically more efficient. Our contributions are as follows.
  • We introduce diffusion resampling, a new resampling paradigm that instantly enables automatic differentiation for $\partial X_{i}^{\theta,*}/\,\partial\theta$, and consequently also for the expectation $\partial\operatorname{\mathbb{E}}\bigl[X_{i}^{\theta,*}\bigr]/\,\partial\theta$. We apply the method for filtering and gradient-based parameter estimation in state-space models with SMC samplers.
  • We prove that our diffusion resampling method is consistent in the number of samples. We show an informative error bound in Wasserstein distance, explicitly quantifying the error propagation of the resampling.
@@ 2 Diffusion differentiable resampling
In this section we show how we can make use of a diffusion model (without training, cf. Baker et al., 2025; Wan and Zhao, 2025) to construct a differentiable resampling scheme and apply it for sequential Monte Carlo. The idea is akin to Equation (2), which computes a linear transportation map, but here we instead use a diffusion model to construct a non-linear map. We define the diffusion model via a (forward-time) Langevin stochastic differential equation (SDE)
@@
initialised at the target $\pi$, where $\pi_{\mathrm{ref}}$ is a user-chosen reference distribution from which we can easily sample (e.g., Gaussian), $W$ is a Brownian motion, and $b$ is a dispersion constant. Under mild conditions (Meyn and Tweedie, 2009), the marginal distribution $p_{t}$ of $X(t)$ converges to $\pi_{\mathrm{ref}}$ geometrically fast as $t\to\infty$. Importantly, Song et al. (2021); Anderson (1982) show that we can leverage this construction to sample from $\pi$ if we can sample $p_{T}$ at some terminal time $T>0$ and simulate the reverse-time SDE [^1] To simplify later analysis we assume that the Brownian motions in Equations (3) and (4) are the same.
@@
At time $T$, the marginal distribution $q_{T}$ of $U(T)$ equals $\pi$ by construction, since X(Tt)X(T-t) and $U(t)$ solve the same Kolmogorov forward equation. Resampling can thus be achieved by sampling from this reversal at $T$. The challenge is that the score function $\operatorname{\nabla\!}\log p_{t}$ is intractable, and in the context of generative diffusion models the score is usually learnt from samples of $\pi$, introducing demanding computations (Zhao et al., 2025). However, given that we have access to $\{(w_{i},X_{i})\}_{i=1}^{N}\sim\pi$, we can approximate the score (Bao et al., 2024) without training according to
@@
where $\alpha_{i}(x,t)\coloneqq w_{i}\,p_{t|0}(x{\;|\;}X_{i})\,/\sum_{j=1}^{N}w_{j}\,p_{t|0}(x{\;|\;}X_{j})$ stands for the normalised weight. This approximation exactly functions as importance sampling, where $\pi(x_{0})$ and $p_{t|0}(\cdot{\;|\;}x_{0})$ stand for the prior/proposal and likelihood, respectively. As such, the established $N\to\infty$ consistency properties of importance sampling apply (Chopin and Papaspiliopoulos, 2020) at least pointwise for $(x,t)\mapsto s_{N}(x,t)$. Evaluation of the function $s$ has an $O(N)$ computational cost if implemented naïvely, and a logarithmic cost if implemented in parallel (Lee et al., 2010).
Therefore, the resampling can be approximately achieved by simulating the reverse SDE (4) using the ensemble score $s_{N}$ in Equation (6) until a terminal time $T$. Similar to Corenflos et al. (2021), this diffusion process too defines an optimal transportation but in the sense of Jordan–Kinderlehrer–Otto scheme (Jordan et al., 1998). The key distinction is that this map is given by construction, and does not need to be computed like in OT with Sinkhorn. Although the diffusion also assumes $T\to\infty$, we show in Section 3 that $T$ scales better than the entropy parameter $1\,/\,\varepsilon$ as a function of $N$.
@@ 2.1 Differentiable sequential Monte Carlo
The diffusion resampling in Algorithm 1 is particularly useful for SMC and SSMs for two reasons. First, the resampling method is pathwise-differentiable by construction. Moreover, recent advances provide methods for propagating gradients through SDE solvers (see, e.g., Bartosh et al., 2025; Li et al., 2020; Kidger et al., 2021) and these methods have been well implemented in common automatic differentiation libraries (Kidger, 2021). Secondly, we can fully leverage the sequential structure of SMC to adaptively choose for a suitable reference distribution $\pi_{\mathrm{ref}}$ based on any previous SMC marginal distribution. To see these aspects, let us begin by considering a parametrised Feynman–Kac model
$Q^{\theta}_{0:J}(z_{0:J})=\frac{1}{L(\theta)}\prod_{j=0}^{J}M_{j}^{\theta}(z_{j}{\;|\;}z_{j-1})\,G_{j}^{\theta}(z_{j},z_{j-1}),$ (8)
where $M^{\theta}_{j}$ and $G_{j}^{\theta}$ are Markov transition and potential functions, respectively, and $L(\theta)$ is the marginal likelihood that we often aim to maximise. Take an SSM with state transition $p_{\theta}(z_{j}{\;|\;}z_{j-1})$ and measurement $p_{\theta}(y_{j}{\;|\;}z_{j})$ for example. In this case a bootstrap construction of the corresponding Feynman–Kac model is simply $M_{j}^{\theta}(z_{j}{\;|\;}z_{j-1})=p_{\theta}(z_{j}{\;|\;}z_{j-1})$ and $G_{j}^{\theta}(z_{j},\cdot)=p_{\theta}(y_{j}{\;|\;}z_{j})$. One can draw samples of a Feynman–Kac model with an SMC sampler as in Algorithm 2. For detailed exposition of Feynman–Kac models and SMC samplers, we refer the readers to Chopin and Papaspiliopoulos (2020); Del Moral (2004).
@@ 16   Weight $w_{j,i}\leftarrow\overline{w}_{j,i}\,/\,L_{j}(\theta)$
This choice of reference $\pi_{\mathrm{ref}}$ is more flexible compared to that of Corenflos et al. (2021) who explicitly use the predictive samples $\{(w_{j-1,i},Z_{j,i})\}_{i=1}^{N}$ at the $j$-th SMC step as the reference to resample $\{(w_{j,i},Z_{j,i})\}_{i=1}^{N}$. In contrast, diffusion resampling allows to use, e.g., the posterior samples $\{(w_{j,i},Z_{j,i})\}_{i=1}^{N}$ to establish the reference, which can be more informative than the predictive one.
@@ 2.2 Mean-reverting Gaussian reference
A remaining question is how to choose the reference distribution $\pi_{\mathrm{ref}}$. In generative sampling, the reference is usually a unit Normal. However, this becomes suboptimal for resampling when the target distribution $\pi$ is geometrically far away from $\mathrm{N}(0,I_{d})$, and as a consequence we would need large enough $T$ for convergence. A more informative choice of $\pi_{\mathrm{ref}}$ is a Gaussian approximation of $\pi$. Given that we have access to the target samples $\{(w_{i},X_{i})\}_{i=1}^{N}\sim\pi$, we can make use of moment matching, where $\mu_{N}\coloneqq\sum_{i=1}^{N}w_{i}\,X_{i}$ and $\Sigma_{N}\coloneqq\sum_{i=1}^{N}w_{i}\,(X_{i}-\mu_{N})\,(X_{i}-\mu_{N})^{\mkern-1.5mu\mathsf{T}}$ stand for the empirical mean and covariance, respectively (Yang et al., 2013; Kang et al., 2025). We can then choose the reference measure to be
@@
$\begin{split}p_{t|0}(x_{t}{\;|\;}x_{0})&=\mathrm{N}(x_{t};m_{t}(x_{0}),V_{t}),\\ m_{t}(x_{0})&\coloneqq x_{0}\,\mathrm{e}^{-b^{2}\,\Sigma_{N}^{-1}\,t}+\mu_{N}\,\bigl(1-\mathrm{e}^{-b^{2}\,\Sigma_{N}^{-1}\,t}\bigr),\\ V_{t}&\coloneqq\Sigma_{N}\,\bigl(1-\mathrm{e}^{-2\,b^{2}\,\Sigma_{N}^{-1}\,t}\bigr).\end{split}$
To combine with the SMC in Algorithm 2, we compute $\mu_{N}$ and $\Sigma_{N}$ based on $\bigl\{(w_{j-1,i},Z_{j-1,i})\bigr\}_{i=1}^{N}$, see Appendix A for details. It is also possible to leverage any Gaussian filter which is commonly used for approximating the optimal proposal (van der Merwe et al., 2000), to establish the reference. The central gist here is to exploit the sequential structure of SMC to adaptively and informatively choose the reference, instead of assuming a static and uninformative one, such as $\mathrm{N}(0,I_{d})$. Similar ideas of using mean-reverting SDEs for informative generative sampling have been used in domain-specific applications, such as image restoration (Luo et al., 2023b, 2025).
In practice, to avoid solving the inversion $\Sigma_{N}^{-1}$, we can approximate it as a diagonal, or directly estimate an empirical precision matrix (Yuan and Lin, 2007; Fan et al., 2016). When the Gaussian construction is insufficient for multi-mode targets, one may use a Gaussian mixture, although the associated semigroup needs to be approximated. Another option is to transform with a diffeomorphism to obtain a flexible yet tractable reference process (Deng et al., 2020).